A boater travels 16 miles per hour on the water on a still day. During one particular windy day, he finds that he travels 48 miles with the wind behind him in the same amount of time that he travels 16 miles into the wind. find the rate of wind.
Answer by checkley77(12844)

D=RT
48=(16+W)T
T=48/(16+W)
16=(16-W)T
T=16/(16-W)
48/(16+W)=16/(16-W) CROSS MULTIPLY.
16(16+W)=48(16-W)
256+16W=768-48W
16W+48W=768-256
64W=512
W=512/64
W=8 MPH. IS THE RATE OF THE WIND.
PROOF:
48/(16+8)=16/(16-8)
48/24=16/8
2=2